OpenStudy (anonymous):

Factor: c2 - d2 Answer (c + d)(c - d) (c + d)(c + d) (c + 2d)(c - d)

OpenStudy (anonymous):

Difference of squares!

OpenStudy (anonymous):

\[ a^2-b^2=(a-b)(a+b) \]

OpenStudy (yttrium):

It's factors are sum and difference of binomials. :)

OpenStudy (anonymous):

Medal

OpenStudy (anonymous):

its A .

OpenStudy (uri):

c(c)-d(d) (c-d)(c+d)

OpenStudy (uri):

or (c+d)(c-d)

OpenStudy (anonymous):

@uri:your answer is true but how can we have (c-d)(c+d) from c(c)-d(d) ? can u prove it ?:)

OpenStudy (anonymous):

Distributive property of multiplication over addition.\[ \huge QED \]

OpenStudy (uri):

Well, c(c)-d(d) = (c+d)(c-d) or (c-d)(c+d)

OpenStudy (anonymous):

Well, c(c)-d(d) = (c+d)(c-d) or (c-d)(c+d)

OpenStudy (anonymous):

Distributive property of multiplication over addition. QED

OpenStudy (anonymous):

@uri : :) I wanted prove not answer again :)

OpenStudy (anonymous):

No problem :) Thank u :)

OpenStudy (anonymous):

\[\begin{split} c(c)-d(d) &=c(c)-d(d)+0 \\ &=c(c)-d(d)+(cd-dc) \\ &=c(c)-d(d)+c(d)-d(c) \\ &=c(c)+c(d)-d(d)-d(c) \\ &=[c(c)+c(d)]-[d(d)+d(c)] \\ &=c(c+d)-d(d+c) \\ &=(c-d)(c+d) \\ \end{split} \]

OpenStudy (anonymous):

@E.ali Medal, now.

OpenStudy (uri):

medal me.kthxbye.

OpenStudy (anonymous):

i dont know who i should give medal too o.O

OpenStudy (anonymous):

To me. Please

OpenStudy (uri):

To me duh.I gave the ansur first.

OpenStudy (uri):

Also,I'm a pretty girl.

OpenStudy (anonymous):

ali gave me first answer :D

OpenStudy (uri):

Whose the prettiest gets the medal.

OpenStudy (anonymous):

i dont see ur picture D:

OpenStudy (uri):

Eli didn't even explain you xD

OpenStudy (anonymous):

ahahahaha :D

OpenStudy (uri):

Be like dat.Uri won't half ya with ansurs again.

OpenStudy (debbieg):

Nobody should get a medal for giving the ANSWER. The medal should go to the person who explained/taught HOW to answer it. (Even if begging for a medal is pretty pathetic.)

OpenStudy (anonymous):

uri, u give to ali then, and i give u :D

OpenStudy (uri):

ANSUR* @DebbieG

OpenStudy (anonymous):

lololol